What is the formula for problems like "Find how many years it will take for Kim's father, who is 39, to be twice as old as Kim, who is 15?" By making a chart I got 9 years in between their current ages and the final age, but I could not come up with an appropriate formula. Best I got was knowing that(n+1)may be included.
Thank you, Jenelle

Answer by rothauserc(4718)

let x represent the years necessary,
39 + x = (15 + x)
39 + x = 2(15 + x)
x = 9 years
check answer
in 9 years Kim's father will be 48 and Kim will be 24